Jack bought 3 items that cost $25 each, and 5 items that cost $40 each. What was the total cost, including 5% sales tax?
Answer by nyc_function(2733)

3 * 25 = 75
5 * 40 = 200
Let t = total cost
t = 0.05(200 + 75) + 275
t = $288.75